OpenStudy (anonymous):

Which values for x and y make the statement (x + 3)(y – 2) = 0 true? x = -3; y = 2 x = -3; y = -2 x = 3; y = 2 x = 3; y = -2 Which of the following gives an example of a set that is closed under multiplication? The product of an even number and an even number The product of a prime number and a prime number. The product of a negative number and a negative number The product of a perfect cube and a perfect cube

OpenStudy (anonymous):

B

OpenStudy (anonymous):

A

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

@BAdhi

OpenStudy (badhi):

if either x = -3 OR y = 2 is there the statement is true

OpenStudy (badhi):

product of two even numbers will give an even number so its closed under multiplication product of two prime numbers will not give a prime number (since the resultant number can be divided from one of the prime numbers that you began with) product of two negetive numbers will not give a negetive number (two negative numbers multiplied will result in a positive number) \(x^3 \times y^3 = (xy)^3\) xy is a perfect cube

Directrix (directrix):

@lauren.ascough http://openstudy.com/code-of-conduct Give Help, Not Answers –I will encourage and guide those needing help, and not just give them an answer